40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4531     Entry time: Fri Apr 15 13:40:00 2011
Author: Larisa Thorne 
Type: Update 
Category: Electronics 
Subject: Low pass filter for X arm laser temperature control, second try 

Plotting the data points yielded by the spec analyzer of my first LPF yielded a result that was not expected: the desired cutoff frequency wasn't achieved because of some extra 100k resistance that wasn't taken into consideration. (see  here ). I have redrawn the Bode graphs for this configuration so that it is easier to see that it is wrong (first attachment)

 

After some calculation adjustments, it was found that the capacitor value could remain at 10uF, but the resistance needed to be changed to 100k to maintain a gain of 0.5 and critical frequency at 0.1Hz. Second attachment is the Bode graph that results from this configuration.

 

Note: Bode graphs are both in Log-Linear scales (Wikipedia said so)

 

Attachment 1: Bode2.jpeg  20 kB  | Hide | Hide all
Bode2.jpeg
Attachment 2: Bode100k.jpg  34 kB  | Hide | Hide all
Bode100k.jpg
ELOG V3.1.3-